A right triangle's height is 3 times the length of its base. If the area of the triangle is 486, what is the height?
Answer by waynest(281)

height = 3x
base = x
Area = 486
formula = 1/2bh
1/2(x)(3x) = 486
1/2(3x^2) = 486
2(1/2)(3x^2) = 2(486)
3x^2 = 972
3x^2/3 = 972/3
x^2 = 324
sqrtx^2 = sqrt324
x = 18
check:
1/2(18)(3(18)) = 486
1/2(18)(54) = 486
1/2(972) = 486
486 = 486